1. Suppose the demand and supply functions for good X are: • Qd = 50 – 10p • Qs = -25 + 15p a) What are the equilibrium price and quantity? b) What is the market outcome if the price is $2.75? What do you expect to happen? Why? c) What is the market outcome if the price is $4.25? What do you expect to happen? Why? d) What happens to equilibrium price and quantity if the demand function becomes Qd = 55 – 5p? e) What happens to the equilibrium price and quantity if the supply function becomes Qs = -40 + 10p (demand is Qd = 50- 10p)?


1
Expert's answer
2020-11-18T14:58:46-0500

a) In equilibrium Qd = Qs, so:

"50 \u2013 10p = -25 + 15p,"

p = $3,

Q = 50 - 10×3 = 20 units.

b) If the price is $2.75, then there will be a shortage, as this price is below equilibrium.

c) If the price is $4.25, then there will be a surplus, as this price is above equilibrium.

d) If the demand function becomes Qd = 55 – 5p, then the demand curve will shift to the right, and the equilibrium price and quantity will increase.

e) If the supply function becomes Qs = -40 + 10p, then the supply curve will shift to the left, the equilibrium price will increase and the equilibrium quantity will decrease.